OpenStudy (anonymous):

If sally can paint a house in 4 hours, and john can paint the same house in 6 hours. How long will it take if they paint the house together?

OpenStudy (anonymous):

1 hour and 20 minutes I think

OpenStudy (anonymous):

I don't remember a formula for this, but imagine a house requiring 100 units of paint. Sally does it in 4 hours, that's 25 units per hour. John does it in 2 hours, or 50 units per. Together, it's 75 units per hour. (100/75) = (4/3) = 1 hr 20 mins

OpenStudy (radar):

Since there is only two, use the product over sum. (6 x 4)/(6+4)

OpenStudy (radar):

Or 24/10=2.4 hours

OpenStudy (amistre64):

How much does each pain tin 1 hour. S = 1/4 of a house J = 1/6 of a house S+J = how much they paint together in 1 hour: 1/4 + 1/6 = (6+4)/24 = 10/24 = 5/12 Together they paint 5/12 part of the house in one hour. x amount of hours times (5/12) will equal 12/12, or the whole thing. (5/12)x = 12/12 = 1 x = 1(12/5) = 12/5 x= 2 and 2 fifths of an hour. what is 2/5 of an hour? Its (60minutes/5) times 2 12(2) = 24 minutes 2 hours and 24 minutes to paint the house.

OpenStudy (anonymous):

Oh, I used 4 hrs and 2 hrs, oops.

OpenStudy (amistre64):

Krb is right if they are unionized :)

OpenStudy (radar):

A hint, when the two work together, the job should be completed in less time of the quickest one.

OpenStudy (amistre64):

before I knew how to work these, I looked at it from my carpentry years and thought.... when i had help it usually slowed me down :)

OpenStudy (radar):

LOL I guess I should say theoretically